# Distinguishing Distance Duality breaking models using electromagnetic and gravitational waves measurements

### v1

@@ Acknowledgments
CDL thanks iniziativa specifica INFN for financial support. This work was partially supported by the research grant number 2022E2J4RK ”PANTHEON: Perspectives in Astroparticle and Neutrino THEory with Old and New messengers” under the program PRIN 2022 funded by the Italian Ministero dell’Università e della Ricerca (MUR). MM acknowledges funding by the Agenzia Spaziale Italiana (asi) under agreement no. 2018-23-HH.0 and support from INFN/Euclid Sezione di Roma. RD acknowledges financial support from INFN – Sezione di Roma 1, esperimento Euclid, and work from COST Action CA21136 – Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by European Cooperation in Science and Technology. CDL and MM acknowledge financial support from Sapienza Università di Roma, provided through Progetti Medi 2021 (Grant No. RM12117A51D5269B). This work made use of Melodie, a computing infrastructure funded by the same project, and PLEIADI, a computing infrastructure installed and managed by INAF.

### v2

@@ Acknowledgments
CDL thanks iniziativa specifica INFN for financial support. This work was partially supported by the research grant number 2022E2J4RK ”PANTHEON: Perspectives in Astroparticle and Neutrino THEory with Old and New messengers” under the program PRIN 2022 funded by the Italian Ministero dell’Università e della Ricerca (MUR). MM acknowledges funding by the Agenzia Spaziale Italiana (asi) under agreement n. 2024-10-HH.0 and support from INFN/Euclid Sezione di Roma. RD acknowledges financial support from INFN – Sezione di Roma 1, esperimento Euclid, and work from COST Action CA21136 – Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse) supported by European Cooperation in Science and Technology. CDL and MM acknowledge financial support from Sapienza Università di Roma, provided through Progetti Medi 2021 (Grant No. RM12117A51D5269B). This work made use of Melodie, a computing infrastructure funded by the same project, and PLEIADI, a computing infrastructure installed and managed by INAF.